if each of the sides of a square is lengthened by 4 m, the area becomes 49 mē. find the length of a side of the original square
Answer by mangeshpjoshi(15)

dear student,
let the side of original square=x metre
so its square =x^2
now if we increase length by 4 m
then new length = (x+4)m
its area = (x+4)^2=49
x^2+8x+16 = 49
x^2+8x-33=0
x^2+11x-3x-33=0
x(x+11)-3(x+11)=0
(x+11)(x-3)=0
from this
x= -11 or x=3
-11= not possible of length so 3 is the correct answer
of oringinal square
thanks
faithfully
mangesh joshi